# Pure minimal injective resolutions and perfect modules for lattices

### v1

@@ Abstract.
In a recent article, Iyama and Marczinzik showed that a lattice is distributive if and only if the incidence algebra is Auslander regular, giving a new connection between homological algebra and lattice theory. In this article we study when a distributive lattice has a pure minimal injective coresolution, a notion first introduced and studied in a work of Ajitabh, Smith and Zhang. We will see that this problem naturally leads to studying when certain antichain modules are perfect modules. We give a classification of perfect antichain modules under the assumption that their canonical antichain resolution is minimal and use this to give a completion classification in lattice theoretic terms of incidence algebras of distributive lattices with pure minimal injective coresolution.
@@ 5. Classification of distributive lattices with perfect injective modules
In forthcoming work the previous problem will be studied for geometric lattices.

### v2

@@ Abstract.
In a recent article, Iyama and Marczinzik showed that a lattice is distributive if and only if the incidence algebra is Auslander regular, giving a new connection between homological algebra and lattice theory. In this article we study when a distributive lattice has a pure minimal injective coresolution, a notion first introduced and studied in a work of Ajitabh, Smith and Zhang. We will see that this problem naturally leads to studying when certain antichain modules are perfect modules. We give a classification of perfect antichain modules under the assumption that their canonical antichain resolution is minimal and use this to give a complete classification in lattice theoretic terms of incidence algebras of distributive lattices with pure minimal injective coresolution. We use our results to answer a question raised by Ajitabh, Smith and Zhang by showing that there exist Auslander-Gorenstein polynomial identity rings without a pure injective coresolution.
@@ 5. Classification of distributive lattices with perfect injective modules
In forthcoming work the previous problem will be studied for geometric lattices. As a final application of our results we answer the following question raised by Ajitabh, Smith and Zhang in the beginning of section 4 in [ASZ2]:
Question 5.7.
Let $A$ be an Auslander-Gorenstein ring that is also a polynomial identity ring. Does $A$ have a pure minimal injective coresolution?
Note here that in [ASZ2] the authors just talk about a pure injective coresolution, but for finite dimensional algebras a pure injective coresolution exists if and only if there is a pure minimal injective coresolution. Here we recall that a ring $R$ is called polynomial identity ring if there is, for some $N>0$, a polynomial $F\neq 0$ in the free algebra, $\mathbb{Z}\langle x_{1},x_{2},\ldots,x_{N}\rangle$, over the ring of integers in N variables $x_{1},x_{2},\ldots,x_{N}$ such that $F(r_{1},r_{2},\ldots,r_{N})=0$ for all $N$-tuples $r_{1},\ldots,r_{N}$ of elements in $R$. Now by the following result of Feinberg every incidence algebra of a finite poset is a polynomial identity ring:
Theorem 5.8.
[Fe, Theorem 2] Let $P$ be a finite poset with incidence algebra $A=KP$ and let $F(x_{1},\ldots,x_{k}):=\sum\limits_{\pi\in S_{k}}{\operatorname{sign}\pi\cdot x_{\pi(1)}\cdot\ldots\cdot x_{\pi(k)}}$ be the standard polynomial in $k$-variables, where $S_{k}$ denotes the symmetric group in $k$ letters and $\operatorname{sign}\pi$ the signum of the permutation $\pi$. Then $A$ is a polynomial identity algebra with $F(r_{1},\ldots,r_{2n})=0$ for all $r_{1},\ldots,r_{2n}\in A$, where $n$ is such that all chains of the poset $P$ have length $\leq n-1$.
Now using our main result Theorem 5.4, we can give a distributive lattice $L$ with incidence algebra $A$ such that $A$ is Auslander-Gorenstein and a polynomial identity ring (by Theorem 5.8) but such that $A$ does not have a pure injective coresolution. The simplest example of such an $A$, where we do not have a pure injective coresolution of $A$ and $A^{op}$ is given in the following example:
Example 5.9.
Let $L$ be the distributive lattice with Hasse diagram
6{6}5{5}3{3}4{4}2{2}1{1}
Let $A$ be the incidence algebra of $L$. Then obviously $A\cong A^{op}$ and by Theorem 5.4 AA and $A^{op}$ do not have a pure injective coresolution. But as $L$ is distributive, AA is Auslander-Gorenstein and by Theorem 5.8 AA is a polynomial identity ring, answering Question 5.7 in the negative.